# The Disparate Benefits of Deep Ensembles

### v1

@@ Abstract
Ensembles of Deep Neural Networks, Deep Ensembles, are widely used as a simple way to boost predictive performance. However, their impact on algorithmic fairness is not well understood yet. Algorithmic fairness investigates how a model’s performance varies across different groups, typically defined by protected attributes such as age, gender, or race. In this work, we investigate the interplay between the performance gains from Deep Ensembles and fairness. Our analysis reveals that they unevenly favor different groups in what we refer to as a disparate benefits effect. We empirically investigate this effect with Deep Ensembles applied to popular facial analysis and medical imaging datasets, where protected group attributes are given and find that it occurs for multiple established group fairness metrics, including statistical parity and equal opportunity. Furthermore, we identify the per-group difference in predictive diversity of ensemble members as the potential cause of the disparate benefits effect. Finally, we evaluate different approaches to reduce unfairness due to the disparate benefits effect. Our findings show that post-processing is an effective method to mitigate this unfairness while preserving the improved performance of Deep Ensembles.
@@ 1 Introduction
Deep Ensembles (Lakshminarayanan et al., 2017) have demonstrated their efficacy as a simple and robust method to improve the performance of individual Deep Neural Networks (DNNs). Their superior performance has made them a popular choice for real-world applications (Bhusal et al., 2021; Dolezal et al., 2022), including high-stakes scenarios where the impact on people’s lives of machine learning supported decisions can be profound, such as in healthcare, education, finance or the law. In such applications, it is crucial to examine how these models perform across different groups that are defined by a protected attribute (e.g., gender, age, race, etc.) which is the focus of the field of Algorithmic Fairness (Barocas et al., 2023). Ensuring equitable operation of these models across protected groups is imperative, as they can significantly impact individuals and communities, potentially widening existing disparities if not adequately addressed. Although the differences in performance across protected groups (group fairness violations) of individual DNNs has been extensively studied (Zhang et al., 2018; Sagawa et al., 2020; Zhang et al., 2022; Arnaiz-Rodriguez & Oliver, 2024), the impact on fairness of ensembling these networks remains underexplored.
In this paper, our aim is to fill this gap by conducting an extensive empirical study of the fairness implications of Deep Ensembles, analyzing their underlying causes, and exploring mitigation strategies. Our empirical study is based on two popular facial analysis datasets and a widely used medical imaging dataset, each with multiple targets and protected group attributes. We evaluate a total of fifteen tasks across five different DNN model architectures and using three standard group fairness measures. Our analyses reveal that Deep Ensembles unevenly benefit different protected groups in what we refer to as the disparate benefits effect (c.f. Fig.1). We further investigate the causes of this disparate benefits effect, and find evidence that differences in the predictive diversity of ensemble members across groups are the reason why ensembling benefits groups differently. Finally, we explore potential approaches to mitigate the negative impact on fairness caused by the disparate benefits effect. We find that Deep Ensembles are more sensitive to the prediction threshold than individual models due to their improved calibration. This makes post-processing methods a suitable approach to mitigate the fairness violations. In fact, our results show that Hardt post-processing (Hardt et al., 2016) is very effective, yielding fairer predictions while preserving the improved performance of Deep Ensembles. In sum, the main contributions of this paper are three-fold:
@@
    We empirically analyze how the performance improvements of Deep Ensembles distribute across groups defined by protected attributes (Sec. 5). Our findings reveal that Deep Ensembles yield disparate benefits across groups, often benefiting the already advantaged group.
@@
    We investigate potential causes for the disparate benefits effect (Sec. 6). Our analysis suggests the per-group differences in the predictive diversity of ensemble members as an underlying factor.
@@
    We evaluate approaches to mitigate the negative impact of the disparate benefits effect (Sec. 7). We find that Deep Ensembles are more sensitive to the prediction threshold due to their improved calibration. Thus, Hardt post-processing (Hardt et al., 2016) is found to be very effective, ensuring more fair predictions while preserving the improved performance of Deep Ensembles.
Figure 1: Negative consequences of the disparate benefits effect of Deep Ensembles. The performance increases, but the fairness decreases when more members are added to the ensemble. Performance is measured by accuracy (FairFace and UTKFace) and AUROC (CheXpert). Fairness is measured as 1-SPD, 1-EOD and 1-AOD, where SPD, EOD and AOD are common metrics capturing group fairness violations. The dashed blue line indicates the average fairness of individual ensemble members. Results for the FairFace and UTKFace datasets are obtained for target age and protected group attribute gender. Results for the CheXpert dataset are obtained for target no finding and protected group attribute age. Statistics are based on five independent runs (ResNet50).
@@ 2 Related Work
Algorithmic Fairness. Algorithmic fairness is defined using various ethical and legal concepts (Barocas & Selbst, 2016; Corbett-Davies et al., 2017; Binns, 2018), resulting in diverse statistical and causal notions of equality across tasks and contexts (Kusner et al., 2017; Mehrabi et al., 2021). We focus on group fairness metrics —statistical discrimination metrics for classification (Carey & Wu, 2023)— that measure error rate differences between groups defined by protected attributes (Hardt et al., 2016; Zafar et al., 2017). Several metrics quantify group fairness by imposing independence conditions on the joint distribution of targets, predictions, and protected attributes (Barocas et al., 2023), capturing performance disparities due to varying input and target distributions among protected groups (Garg et al., 2020; Pombal et al., 2022). Consequently, a multitude of ML techniques have emerged over the past decade to promote group algorithmic fairness (Mehrabi et al., 2021) by modifying the data (pre-processing) (Kamiran & Calders, 2012; Arnaiz-Rodriguez & Oliver, 2024), the learning process (in-processing) (Agarwal et al., 2018; Jung et al., 2023); or the model’s decision rule (post-processing) (Hardt et al., 2016; Cruz & Hardt, 2024). In this paper, we focus on group algorithmic fairness and analyze the impact on group fairness of Deep Ensembles.
@@
Deep Ensembles. Deep Ensembles (Lakshminarayanan et al., 2017) are known as a simple and effective method to boost the performance of DNNs and to estimate uncertainty (Ovadia et al., 2019; Ashukha et al., 2020; Schweighofer et al., 2023). They mostly rely on the stochasticity of the initialization and optimization procedure for diversity (Fort et al., 2019). However, obtaining more diverse Deep Ensembles is still an active area of research (Rame & Cord, 2021; Lee et al., 2023; Pagliardini et al., 2023). Furthermore, the exact mechanisms that yield the performance improvements observed in Deep Ensembles remain an open research question (Abe et al., 2022b; Jeffares et al., 2023; Abe et al., 2024). Prior work at the intersection of algorithmic fairness and ensembling has investigated the effect of model multiplicity (Marx et al., 2020; Coston et al., 2021; Black et al., 2022a, b; Long et al., 2023; Cooper et al., 2024), and has reported that ensembling decreases the multiplicity of predictions, thus being less arbitrary than individual models. Shallow model ensembles have been used to improve the fairness of outcomes (Kamiran & Calders, 2012), yet we are not aware of any work that has investigated the impact of Deep Ensembles on group fairness metrics.
@@
The most closely related previous work to ours is Ko et al. (2023), which investigates the effect of Deep Ensembles on subgroup performance and served as an inspiration to our work. However, their focus and methodology are different from ours. For most of their experiments, the group variable of interest $A$ is defined as a subset of the full target space $\mathcal{Y}$, i.e., of the worst and best performing targets. In our experiments with real-world data, groups are defined by the values of a protected attribute, such as age, gender, or race. Furthermore, Ko et al. do not consider established group fairness measures as we do, focusing instead on per-group changes in accuracy. Finally, Ko et al. conclude that Deep Ensembles have exclusively positive impact, while we show that they can negatively affect group fairness. We investigate potential causes for this effect and analyze mitigation strategies that preserve fairness while maintaining the performance gains of the ensembles.
@@ 3 Background
We consider the canonical setting of binary classification with inputs $\bm{x}\in\mathbb{R}^{D}$, targets $y\in\{0,1\}$, and group attributes $a\in\{0,1\}$ defined according to protected or sensitive variables, such as gender, age, or race. Furthermore, we consider DNNs as the models to map an input $\bm{x}$ to the 1-dimensional probability simplex $\Delta^{1}=\left\{(s_{0},s_{1})\in\mathbb{R}^{2}\mid s_{0}\geq 0,s_{1}\geq 0,s_{0}+s_{1}=1\right\}$. We define this mapping as $f_{\bm{w}}:\mathbb{R}^{D}\rightarrow\Delta^{1}$ for a model with parameters $\bm{w}$. The output of this mapping defines the distribution parameters of the predictive distribution of the model, denoted by $p(y\mid\bm{x},\bm{w})$. A training dataset $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{J}$ is used to determine the model parameters by minimizing the cross-entropy loss. The final prediction $\hat{y}$ is given by the argmax over the predictive distribution.
Deep Ensembles. Deep Ensembles (Lakshminarayanan et al., 2017) are an ensemble method that uses DNNs as the base learners. While shallow learners aggregate ensemble predictions via majority voting, Deep Ensembles typically average the output distributions of individual members. Furthermore, individual models are generally trained independently on the same data using different random seeds for initialisation and training. Deep Ensembles are widely recognized as a way to perform approximate sampling from the posterior distribution $p(\bm{w}\mid\mathcal{D})=p(\mathcal{D}\mid\bm{w})p(\bm{w})/p(\mathcal{D})$ (Wilson & Izmailov, 2020; Ashukha et al., 2020), often providing the most faithful posterior approximations (Izmailov et al., 2021). The predictive distribution of an ensemble with $N$ members is given by
$p(y\mid\bm{x},\mathcal{D})\ =\ \int_{W}p(y\mid\bm{x},\bm{w})\ p(\bm{w}\mid\mathcal{D})\ \mathrm{d}\bm{w}\ \approx\ \frac{1}{N}\sum_{n=1}^{N}p(y\mid\bm{x},\bm{w}_{n})\ ,$ (1)
where $\bm{w}_{n}\sim p(\bm{w}\mid\mathcal{D})$. Thus, it is an approximation of the posterior predictive distribution. The prediction of the Deep Ensemble, equivalent to a single model, is given by $\hat{y}=\mathop{\mathrm{argmax}\,}p(y\mid\bm{x},\mathcal{D})$.
Group Fairness. Group fairness desiderata are based on the statistical dependencies between the random variables of the predicted outcomes $\hat{Y}$, the observed outcomes $Y$ and the protected group attribute $A$. Barocas et al. (2023) define three conditions to evaluate algorithmic fairness: independence ($\hat{Y}\bot A$), separation ($\hat{Y}\bot A\mid Y$), and sufficiency ($Y\bot A\mid\hat{Y}$). Following widespread convention, we consider binary outcomes and protected groups, with $\hat{Y}=Y=1$ to be the positive outcome and $A=1$ to be the advantaged group. We measure three different notions of fairness that are common in the literature (Mehrabi et al., 2021; Caton & Haas, 2023).
@@
$P(\hat{Y}=1\mid A=1,Y=1)\ =\ P(\hat{Y}=1\mid A=0,Y=1)\ .$ (3)
Third, equalized odds (Hardt et al., 2016), which is a stricter version of equal opportunity where the predictive independence must hold conditioned on both positive and negative observed outcomes. Equalized odds is thus formally defined as
$P(\hat{Y}=1\mid A=1,Y=y)\ =\ P(\hat{Y}=1\mid A=0,Y=y)\ ,\quad y\in\{0,1\}\ .$ (4)
All operationalized notions of fairness have limitations such that it is not necessarily guaranteed that changing the model predictions to satisfy the conditions given by Eq. (2) - (4) will actually lead to perfectly fair outcomes in the real world (Selbst et al., 2019; Liu et al., 2018). Furthermore, some notions of fairness can be incompatible with each other, such as statistical parity and equalized odds if $A$ and $Y$ are not independent (Chouldechova, 2017; Kleinberg et al., 2017). However, these metrics are still a meaningful and widely used tool to quantify group fairness.
@@ 4 Experimental Setup
Datasets. In our experiments, we evaluated Deep Ensembles on three different vision datasets. First, two facial analysis datasets, namely FairFace (Karkkainen & Joo, 2021) and UTKFace (Zhang et al., 2017). For those datasets, all models were trained on the training split of FairFace and evaluated on the official test split of FairFace and the full UTKFace dataset. Protected group attributes were binarized, except for gender which was already binary. For the attribute age, we defined young and old, where a person is considered old from 40 onwards to obtain a roughly balanced age distribution. For the attribute race, we binarized it into white vs non-white. We trained the models using one of the attributes as target variable and evaluating it with the remaining two attributes as protected group variables for all possible combinations of target and protected group attributes. Second, the CheXpert medical imaging dataset (Irvin et al., 2019) using the recommended targets provided by Jain et al. (2021) and protected group attributes provided by Gichoya et al. (2022). The no finding target was used to train and evaluate the models. Samples without all protected group attributes have been removed. A random subset of 1/8 was split as test dataset. Protected group attributes were binarized as for the facial analysis datasets. Additional details are given in Sec. D.1 in the appendix.
@@
Models and training. We used five different DNN architectures, namely ResNet18/34/50 (He et al., 2016), RegNet-Y 800MF (Radosavovic et al., 2020) and EfficientNetV2-S (Tan & Le, 2021) for our evaluation, due to their widespread adoption and competitive performance in vision tasks. The models that were trained on the FairFace training dataset were trained for 100 epochs using SGD with momentum of 0.9 with a batch size of 256 and learning rate of 1e-2. Furthermore, a standard combination of linear (from factor 1 to 0.1) and cosine annealing schedulers was used. The models that were trained on the CheXpert training dataset were trained for 30 epochs given that the training dataset is roughly thrice the size of FairFace, resulting in a similar number of gradient steps and similar learning rate schedule. We independently trained 10 models for 5 architectures on 4 target variables with 5 seeds. Thus, a total of 1,000 individual models were obtained for our evaluation. The results discussed in the main paper correspond to using ResNet50 as the model architecture. Additional results for other model architectures are provided in Sec. F.1 and Sec. F.2 in the appendix.
@@
Performance Metrics. We utilized accuracy as the performance metric on the FairFace and UTKFace datasets. In the case of CheXpert, we measured performance using the AUROC as established by previous work on this dataset (Zhang et al., 2022; Zong et al., 2023).
@@
Group Fairness Metrics. We measured group fairness using empirical estimators for the fairness desiderata given by Eq. (2) - (4). Statistical Parity Difference (SPD) estimates the violation of the condition given by Eq. (2) and it is computed as
$\text{SPD}\ =\ \text{PR}_{A=1}\ -\ \text{PR}_{A=0}\ ,$ (5)
@@
where $\text{PR}_{A=a}$ is the positive rate calculated on the partition of the test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k},a_{k})\}_{k=1}^{K}$ with the corresponding protected group attribute $a$. Equal Opportunity Difference (EOD) estimates the violation of the condition given by Eq. (3) and it is expressed as
$\text{EOD}\ =\ \text{TPR}_{A=1}\ -\ \text{TPR}_{A=0}\ ,$ (6)
@@
where $\text{TPR}_{A=a}$ is the true positive rate, calculated for the respective group partitions of the test dataset. Average Odds Difference (AOD) (Bellamy et al., 2018) is an estimator of a relaxation of equalized odds (c.f. Eq. (4)). AOD is computed by
$\text{AOD}\ =\ \frac{1}{2}\left|\text{TPR}_{A=1}\ -\ \text{TPR}_{A=0}\right|\ +\ \frac{1}{2}\left|\text{FPR}_{A=1}\ -\ \text{FPR}_{A=0}\right|\ ,$ (7)
@@
where $\text{FPR}_{A=a}$ is the false positive rate, calculated for the respective group partitions of the test dataset. Due to our assumption that $A=1$ is the advantaged group, all measures are consequently $\in[0,1]$, where 00 is the most fair. More details on Eq. (5) - (7) are given in Sec. A in the appendix.
5 The Disparate Benefits Effect of Deep Ensembles
In this section, we study the disparate benefits effect for Deep Ensembles using the experimental setup described in Sec. 4. First, we investigate the disparate benefits effect on the FairFace (FF) test dataset. Second, we apply the same models trained on FF to the UTKFace (UTK) dataset. UTK contains similar facial images as FF but from a different source, representing a realistic setting for facial analysis under slight distribution shifts. Third, we investigate the disparate benefits effect on the CheXpert (CX) medical imaging dataset to assess whether the impact on fairness of Deep Ensembles also occurs in other domains than facial analysis. Our analysis examines two primary facets of the disparate benefits effect: (i) the relationship between the number of ensemble members and the changes in performance and fairness violations (Fig. 1); and (ii) the targets and protected group attributes where a statistically significant disparate benefits effect is observed (Tab. 1).
Facial analysis (FF). The top row of Fig. 1 shows results for FF, where models were trained on target age and evaluated under the protected group attribute gender. We find that performance increases while fairness decreases when adding ensemble members. In particular, the largest decrease in fairness occurs when the first member is added to the Deep Ensemble. Tab. 1 lists the change ($\Delta$) in performance and fairness violations between the individual models and a Deep Ensemble of 10 members for all tasks. The performance always increases for the Deep Ensemble (positive $\Delta$). However, fairness does not necessarily increase after ensembling. We observe a disparate benefits effect with significant changes in the fairness metrics for four out of six target / protected group combinations. It occurs primarily when individual members already exhibit substantial levels of fairness violations (gray cells in Tab. 1). The strongest disparate benefits effect (largest absolute $\Delta$) has negative impact, thus increasing the fairness violations. However, there are also cases where the Deep Ensemble is a more fair classifier than individual models (negative $\Delta$). Overall, our results show that Deep Ensembles can decrease fairness, necessitating mitigation strategies.
Facial analysis under a distribution shift (UTK). The middle row of Fig. 1 depicts the results on the UTK dataset, with the same target and protected group as for FF (top row). Individual ensemble members exhibit higher fairness violations than for FF, which can be explained by the distribution shift between FF and UTK. However, the magnitude and behavior of the disparate benefits effect when adding ensemble members are similar to those observed with the FF dataset. The results for all target / group combinations are listed in Tab. 1. Findings for UTK are overall similar to those reported on the FF dataset. An notable exception is that the difference in SPD with target variable race and protected group attribute age is of opposite sign and larger for UTK than for FF.
Medical imaging (CX). The bottom row of Fig. 1 shows the results on the CX dataset with age as protected group attribute. The disparate benefits effect also occurs in this task, but with a smaller magnitude, which is explained by the smaller performance gains of Deep Ensembles on this dataset. Similarly as with the facial dataset, the change in fairness after adding the first ensemble member is the most pronounced in this dataset. The complete results for all protected groups are listed in Tab. 1. For the protected group age, the disparate benefits effect occurs under all fairness measures. Moreover, there is a significant difference in SPD for the protected group race, although individual models do not have substantial SPD and vice versa for EOD.
Additional results. We investigate the influence of the model size of the individual ensemble members in Sec. F.1 in the appendix. Our results show that for tasks where the disparate benefits effect occurs, it increases with model size. Furthermore, we also analyze the disparate benefits effect under different model architectures. Results and more details are given in Sec. F.2 in the appendix, finding that the results provided in the main paper are consistent across architectures. Finally, we also show the disparate benefits effect for heterogeneous Deep Ensembles in Sec. F.3 of the appendix.

### v2

@@ Abstract
Ensembles of Deep Neural Networks, Deep Ensembles, are widely used as a simple way to boost predictive performance. However, their impact on algorithmic fairness is not well understood yet. Algorithmic fairness examines how a model’s performance varies across socially relevant groups defined by protected attributes such as age, gender, or race. In this work, we explore the interplay between the performance gains from Deep Ensembles and fairness. Our analysis reveals that they unevenly favor different groups, a phenomenon that we term the disparate benefits effect. We empirically investigate this effect using popular facial analysis and medical imaging datasets with protected group attributes and find that it affects multiple established group fairness metrics, including statistical parity and equal opportunity. Furthermore, we identify that the per-group differences in predictive diversity of ensemble members can explain this effect. Finally, we demonstrate that the classical Hardt post-processing method is particularly effective at mitigating the disparate benefits effect of Deep Ensembles by leveraging their better-calibrated predictive distributions.
Machine Learning, ICML
@@ 1 Introduction
Deep Ensembles (Lakshminarayanan et al., 2017) have demonstrated their efficacy as a simple and robust method to improve the performance of individual Deep Neural Networks (DNNs). Their superior performance has made them a popular choice for real-world applications (Bhusal et al., 2021; Dolezal et al., 2022), including high-stakes scenarios where the impact on people’s lives of Machine Learning (ML) supported decisions can be profound, such as in healthcare, education, finance or the law. In such applications, it is crucial to examine how these models perform across different groups that are defined by a protected attribute (e.g., gender, age, race, etc.) which is the focus of the field of Algorithmic Fairness (Barocas et al., 2023). Ensuring equitable operation of these models across protected groups is imperative, as they can significantly impact individuals and communities, potentially widening existing disparities if not adequately addressed. Although the differences in performance across protected groups (group fairness violations) of individual DNNs has been thoroughly studied (Zhang et al., 2018; Sagawa et al., 2020; Zhang et al., 2022; Arnaiz-Rodriguez & Oliver, 2024), the impact on fairness due to ensembling these networks remains underexplored.
Figure 1: Unfairness caused by the disparate benefits effect of Deep Ensembles. Adding members to the ensemble increases its performance (Accuracy, AUROC) but decreases its fairness (1-SPD, 1-EOD, 1-AOD). We find evidence for this effect across multiple standard fairness metrics (cols) and vision datasets (rows).
In this paper, our aim is to fill this gap by conducting an extensive empirical study of the fairness implications of Deep Ensembles, analyzing their underlying causes, and exploring mitigation strategies. Our empirical study is based on two popular facial analysis datasets and a widely used medical imaging dataset, each with multiple binary target variables and protected group attributes. We evaluate a total of fifteen tasks across five different DNN model architectures and three standard group fairness metrics. Our analyses reveal that Deep Ensembles unevenly benefit different protected groups in what we refer to as the disparate benefits effect (cf. Fig.1). We further investigate the causes of this disparate benefits effect and find evidence that differences in the predictive diversity of ensemble members across groups can explain why ensembling benefits groups differently. Finally, we explore potential approaches to mitigate the negative impact on fairness caused by the disparate benefits effect. We find that Deep Ensembles are more sensitive to the prediction threshold than individual models due to their improved calibration. This makes post-processing methods a suitable approach to mitigate the fairness violations. In fact, our results show that Hardt post-processing (Hardt et al., 2016) is very effective, resulting in fairer predictions while preserving the improved performance of Deep Ensembles. In sum, the main contributions of this paper are threefold:
@@
    We empirically analyze how the performance gains of Deep Ensembles are distributed across groups defined by protected attributes (Sec. 5). Our findings reveal that Deep Ensembles yield disparate benefits across groups, often benefiting the already advantaged group.
@@
    We investigate the potential causes for the disparate benefits effect (Sec. 6). Our analysis identifies per-group differences in the predictive diversity of ensemble members as a key contributing factor.
@@
    We evaluate approaches to mitigate unfairness due to the disparate benefits effect (Sec. 7). We find that Deep Ensembles are more sensitive to the prediction threshold due to their improved calibration. Thus, Hardt post-processing (Hardt et al., 2016) is found to be very effective, ensuring more fair predictions while preserving the improved performance of Deep Ensembles.
@@ 2 Related Work
Algorithmic Fairness. Algorithmic fairness is defined by means of various ethical and legal concepts (Barocas & Selbst, 2016; Corbett-Davies et al., 2017; Binns, 2018), resulting in diverse statistical and causal notions of equality between tasks and contexts (Kusner et al., 2017; Mehrabi et al., 2021). We focus on group fairness metrics —statistical discrimination metrics for classification (Carey & Wu, 2023)— that measure error rate differences between groups defined by protected attributes (Hardt et al., 2016; Zafar et al., 2017). Several metrics quantify group fairness by imposing independence conditions on the joint distribution of targets, predictions, and protected attributes (Barocas et al., 2023), capturing performance disparities due to varying input and target distributions among protected groups (Garg et al., 2020; Pombal et al., 2022). Consequently, a multitude of ML techniques have emerged over the past decade to promote group algorithmic fairness (Mehrabi et al., 2021) by modifying the data (pre-processing) (Kamiran & Calders, 2012; Arnaiz-Rodriguez & Oliver, 2024), the learning process (in-processing) (Agarwal et al., 2018; Jung et al., 2023); or the model’s decision rule (post-processing) (Hardt et al., 2016; Cruz & Hardt, 2024). In this paper, we focus on group algorithmic fairness and analyze the impact of Deep Ensembles on group fairness.
@@
Deep Ensembles. Deep Ensembles (Lakshminarayanan et al., 2017) are known as a simple and effective method to boost the performance of DNNs and to estimate predictive uncertainty (Ovadia et al., 2019; Ashukha et al., 2020; Schweighofer et al., 2023). They mainly rely on the stochasticity of the initialization and optimization procedure for diversity (Fort et al., 2019). However, obtaining more diverse Deep Ensembles is still an active area of research (Rame & Cord, 2021; Lee et al., 2023; Pagliardini et al., 2023). Furthermore, the exact mechanisms that produce the performance improvements observed in Deep Ensembles remain an open research question (Abe et al., 2022b; Jeffares et al., 2023; Abe et al., 2024).
Ensemble Fairness. Prior work at the intersection of algorithmic fairness and ensembling has investigated the effect of model multiplicity (Marx et al., 2020; Coston et al., 2021; Black et al., 2022a, b; Long et al., 2023; Cooper et al., 2024), and has reported that ensembling decreases the multiplicity of predictions, thus being less arbitrary than individual models. Shallow model ensembles (i.e., models that are not DNNs) have been used to improve the fairness of outcomes (Kamiran & Calders, 2012). Kenfack et al. (2021) considered a fairness-aware weighting of shallow model ensembles. Similarly, Gohar et al. (2023) investigates multiple research questions regarding ensembles of shallow models. Theoretical considerations on ensemble fairness have been investigated by Grgić-Hlača et al. (2017). Based on these insights, Bhaskaruni et al. (2019) proposed a fair ensembling strategy adopting the AdaBoost framework. However, we are not aware of any work that has investigated the impact of Deep Ensembles on group fairness metrics.
@@
The most closely related previous work to ours is that by Ko et al. (2023), which investigates the effect of Deep Ensembles on subgroup performance and served as an inspiration for our work. However, their focus and methodology are different from ours. For most of their experiments, the group variable of interest $A$ is defined as a subset of the full target space $\mathcal{Y}$, i.e., of the worst and best performing targets. In our experiments with real-world data, groups are defined by the values of a protected attribute, such as age, gender, or race. Furthermore, Ko et al. do not consider established group fairness metrics as we do, focusing instead on per-group changes in accuracy. Finally, Ko et al. conclude that Deep Ensembles have exclusively positive impact, while we show that they can negatively affect group fairness. In addition, we investigate potential causes for this effect and analyze mitigation strategies that preserve fairness while maintaining the performance gains of the ensembles.
@@ 3 Background
We consider the canonical setting of binary classification with inputs $\bm{x}\in\mathbb{R}^{D}$, targets $y\in\{0,1\}$, and group attributes $a\in\{0,1\}$ defined according to protected or sensitive variables, such as gender, age, or race. We further consider DNNs as the models to map an input $\bm{x}$ to the 1-dimensional probability simplex $\Delta^{1}=\left\{(s_{0},s_{1})\in\mathbb{R}^{2}\mid s_{0}\geq 0,s_{1}\geq 0,s_{0}+s_{1}=1\right\}$. We define this mapping as $f_{\bm{w}}:\mathbb{R}^{D}\rightarrow\Delta^{1}$ for a model with parameters $\bm{w}$. The output of this mapping defines the distribution parameters of the predictive distribution of the model, denoted by $p(y\mid\bm{x},\bm{w})$. A training dataset $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{J}$ is used to determine the model parameters by minimizing the cross-entropy loss. The final prediction $\hat{y}$ is given by the argmax over the predictive distribution.
Deep Ensembles. Deep Ensembles (Lakshminarayanan et al., 2017) are an ensemble method that uses DNNs as the base learners. While shallow learners often aggregate predictions in the ensemble via majority voting, Deep Ensembles typically average the output distributions of individual members. Furthermore, individual models are generally trained independently on the same data using different random seeds for initialisation and training. Deep Ensembles are widely recognized as a way to perform approximate sampling from the posterior distribution $p(\bm{w}\mid\mathcal{D})=p(\mathcal{D}\mid\bm{w})p(\bm{w})/p(\mathcal{D})$ (Wilson & Izmailov, 2020; Ashukha et al., 2020), often providing the most faithful posterior approximations (Izmailov et al., 2021). The predictive distribution of an ensemble with $N$ members is given by
$p(y\mid\bm{x},\mathcal{D})$ $=\ \int_{W}p(y\mid\bm{x},\bm{w})\ p(\bm{w}\mid\mathcal{D})\ \mathrm{d}\bm{w}$ (1)
$\approx\ \frac{1}{N}\sum_{n=1}^{N}p(y\mid\bm{x},\bm{w}_{n})\ ,\quad\bm{w}_{n}\sim p(\bm{w}\mid\mathcal{D})$
Thus, it is an approximation of the posterior predictive distribution. The prediction of the Deep Ensemble equivalent to a single model is given by $\hat{y}=\mathop{\mathrm{argmax}\,}p(y\mid\bm{x},\mathcal{D})$.
Group Fairness. The group fairness desiderata are based on the statistical dependencies between the random variables of the predicted outcomes $\hat{Y}$, the observed outcomes $Y$ and the protected group attribute $A$. Following widespread convention, we consider binary outcomes and protected groups, with $\hat{Y}=Y=1$ being the positive outcome and $A=1$ the advantaged group. We focus on three well-established notions of group fairness as follows (Mehrabi et al., 2021; Caton & Haas, 2023).
@@
$P(\hat{Y}\!=\!1\mid A\!=\!1,Y\!=\!1)\ =\ P(\hat{Y}\!=\!1\mid A\!=\!0,Y\!=\!1)$ (3)
Third, equalized odds (Hardt et al., 2016), which is a stricter version of equal opportunity where the predictive independence must hold conditioned on both positive and negative observed outcomes. Equalized odds is formally defined as
$P(\hat{Y}\!=\!1\mid A\!=\!1,Y\!=\!y)\ =\ P(\hat{Y}\!=\!1\mid A\!=\!0,Y\!=\!y)$ (4)
$\forall y\in\{0,1\}$. These measures are particularly relevant because they operationalize antidiscrimination principles, such as disparate impact in U.S. law (Feldman et al., 2015). Statistical parity focuses on ensuring equal outcomes, while equal opportunity and equalized odds balance error rates to promote equity across groups. All operationalized notions of fairness have limitations such that it is not necessarily guaranteed that changing the model predictions to satisfy these conditions will actually lead to perfectly fair outcomes in the real world (Selbst et al., 2019; Liu et al., 2018). Furthermore, some notions of fairness can be incompatible with each other, such as statistical parity and equalized odds if $A$ and $Y$ are not independent (Chouldechova, 2017; Kleinberg et al., 2017). Nevertheless, these metrics are a meaningful and widely used tool to quantify group fairness.
@@ 4 Experimental Setup
Datasets. In our experiments, we evaluated Deep Ensembles on three different vision datasets. First, two facial analysis datasets, namely FairFace (Karkkainen & Joo, 2021) and UTKFace (Zhang et al., 2017). For these datasets, all models were trained on the training split of FairFace and evaluated on the official test split of FairFace and the full UTKFace dataset. Protected group attributes were binarized, except for gender which was already binary. For the attribute age, we defined young and old, where a person is considered old from 40 years onwards to obtain a roughly balanced age distribution. For the attribute race, we binarized into white vs. non-white. We trained the models using one of the attributes as the target variable and evaluated with the remaining two attributes as protected group variables for all possible combinations of the target and protected group attributes. Second, the CheXpert medical imaging dataset (Irvin et al., 2019) using the recommended targets provided by Jain et al. (2021) and protected group attributes provided by Gichoya et al. (2022). The no finding target was used to train and evaluate the models. Samples without all protected group attributes have been removed. A random subset of 1/8 was split as test dataset. Protected group attributes were binarized as for the facial analysis datasets. Additional details are given in Apx. D.1.
@@
Models and training. We used five different DNN architectures, namely ResNet18/34/50 (He et al., 2016), RegNet-Y 800MF (Radosavovic et al., 2020) and EfficientNetV2-S (Tan & Le, 2021) for our evaluation, due to their widespread adoption and competitive performance in vision tasks. The models that were trained on the FairFace training dataset were trained for 100 epochs using SGD with momentum of 0.9 with a batch size of 256 and learning rate of 1e-2. Furthermore, a standard combination of linear (from factor 1 to 0.1) and cosine annealing schedulers was used. The models that were trained on the CheXpert training dataset were trained for 30 epochs given that the training dataset is roughly three times the size of FairFace, resulting in a similar number of gradient steps and a similar learning rate schedule. We independently trained 10 models for 5 architectures on 4 target variables with 5 seeds. Thus, a total of 1,000 individual models were obtained for our evaluation. The results discussed in the main paper correspond to the use of ResNet50 as the model architecture. Additional results for the other model architectures are provided in Apx. F.2 and Apx. F.3.
@@
Performance Metrics. We utilized accuracy as the performance metric on the FairFace and UTKFace datasets. In the case of CheXpert, we measured performance using the AUROC as established in previous work on this dataset (Zhang et al., 2022; Zong et al., 2023).
@@
Group Fairness Metrics. We measured group fairness using empirical estimators for the fairness desiderata given by Eq. (2) - (4). Statistical Parity Difference (SPD) estimates the violation of the condition given by Eq. (2) and is computed as
$\text{PR}_{A=1}\ -\ \text{PR}_{A=0}\ ,$ (5)
@@
where $\text{PR}_{A=a}$ is the positive rate calculated on the partition of the test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k},a_{k})\}_{k=1}^{K}$ with the corresponding protected group attribute $a$. Equal Opportunity Difference (EOD) estimates the violation of the condition given by Eq. (3) and it is computed as
$\text{TPR}_{A=1}\ -\ \text{TPR}_{A=0}\ ,$ (6)
@@
where $\text{TPR}_{A=a}$ is the true positive rate, calculated for the respective group partitions of the test dataset. Average Odds Difference (AOD) (Bellamy et al., 2018) is an estimator of a relaxation of equalized odds (cf. Eq. (4)) computed as
$\frac{1}{2}\left|\text{TPR}_{A=1}-\text{TPR}_{A=0}\right|+\frac{1}{2}\left|\text{FPR}_{A=1}-\text{FPR}_{A=0}\right|\ ,$ (7)
@@
where $\text{FPR}_{A=a}$ is the false positive rate, calculated for the respective group partitions of the test dataset. Due to our assumption that $A=1$ is the advantaged group, all measures are consequently $\in[0,1]$, where 00 is the most fair. More details on Eq. (5) - (7) are given in Apx. A.
5 The Disparate Benefits of Deep Ensembles